Suppose we observe the 3-year Treasury security rate (1R3) to be 6 percent, the expected 1-year rate next year—E(2r1)—to be 4 percent, and the expected 1-year rate the following year—E(3r1)—to be 5 percent. If the unbiased expectations theory of the term structure of interest rates holds, what is the 1-year Treasury security rate, 1R1? (Round your answer to 2 decimal places.)
Answer #1

Unbiased expectations states that

(1+1R1)((1+2R1)*(1+3R1)=(1+1R3)^3

=>1R1=1.06^3/(1.04*1.05)-1

=>1R1=9.0674%